# High-order Fusion Graph Contrastive Learning for Recommendation

### v2

@@ 3. HFGCL: HIGH-ORDER FUSION GRAPH CONTRASTIVE LEARNING
In this section, we propose a High-order Fusion Graph Contrastive Learning (HFGCL) framework in Fig. 3. HFGCL generates high quality contrastive views exclusively from high-order information, and efficiently fusing the diverse self-supervised signals by aggregating information from both positive pairs and negative samples.
Figure 3. Framework of HFGCL. Left: User-item interaction. Center: High-order contrastive views generation process. In the figure, let $h=2$ and $L=3$. After aggregating high-order embeddings, user $u_{1}$ has a representation more similar to the positive item $i_{2}$, with only $u_{4}$ missing. This indicates a close relationship between $u_{1}$ and $i_{2}$ in the high-order embedding space. Right: Positive and negative pair sampling approaches for HFGCL, refer to Fig. 4 for the meaning.
3.1. Different Perspectives for Contrastive Views
Following the analysis in Section 2.3, we shift our attention to the traditional recommendation task. Inspired by the natural language processing (NLP) (Gao et al., 2022b), which generates contrastive views by applying the dropout function to input sentence. We believe that the contrastive views same exist in traditional recommendation task.
Considering a different perspective, we directly treat users and items as contrasting views. Given that users and items are first-order neighbors of each other, their high-order information are essentially identical. As graph convolution progresses, users and items gradually capture neighborhood information, resulting in their embeddings becoming progressively closer. Therefore, GCN process ultimately generates a set of distinctive contrastive views.
The previous contrastive views originate from the same sample ($e.g.,$ $\mathbf{e}_{u}^{\prime}$ and $\mathbf{e}_{u}^{\prime\prime}$ from $\mathbf{e}_{u}$), necessitating that a set of views can be selected as negative samples ($e.g.,$ $\sum_{i\in B}\mathbf{e}_{u_{i}}^{\prime\prime}$) to calculate CL losses.
@@ 3.2. High-order Contrastive Views
Many studies use aggregation methods to combine layer-wise embeddings into the final embeddings. Though this method effectively fuses information from both the sample itself and its neighborhoods, we find that the low-order information is extremely unique. This uniqueness can seriously affect the similarity between positive pairs, thereby interfering with the mutual information of them.
@@
3.3. More Negative Samples to CL Loss
In recommender systems, data augmentation typically involve augmenting the original user or item embeddings with multiple GCN operations to generate different contrastive views. When calculating CL loss, different views of the same sample are regarded as positive pairs, while a set of augmented views are selected as negative samples. Among them, it is not difficult to find that the negative samples used in the previous CL loss come from a set of augmented views. In contrast, for the high-order contrastive views propose in Section 3.2, we give the corresponding CL loss in Eq. 4. It is differently from the previous negative samples sampling approach, the negative samples come from the opposing views. This difference leads to the self-supervised signals of users and items themselves are missing.
Original CL Loss. To be able to obtain this part of the self-supervised signal, we attempt to treat the samples themselves as negative samples and introduce the corresponding CL loss:
@@ 3.3. More Negative Samples to CL Loss
where $a$ and $b$ denote different users/items samples. Unlike traditional CL loss, the positive pairs ($i.e.,$ $a$-$a$) of samples are identical. Subsequently, we combine the CL loss associated with the users and the items themselves:
@@
After obtaining the CL objectives for both users and items themselves, we integrate these with our main CL loss:
@@
Fusion CL Loss. Most studies obtain self-supervised signals through traditional CL objectives (e.g., separately obtaining signals from users and items). Although effective, we argue that differences in these self-supervised signals lead to suboptimal recommendations. Instead, in our CL objective, we propose four CL objectives to capture self-supervised signals. While proposed CL objectives capture more signals, in fact, different self-supervised signals make them extremely difficult to effectively integrate these self-supervised signals. It motivates us to further improve our CL loss to fuse these diverse self-supervised signals. Considering that, unlike the traditional CL loss, our proposed CL loss allows both users and items to serve as negative samples. One simple method involves directly concatenating the two contrastive views and using them as negative samples in CL:
(10) $\mathcal{C}_{con}(\mathbf{e}_{u},\mathbf{e}_{i})$ $=-\sum_{<u,i>\in\mathcal{B}}\mathrm{log}\frac{\mathrm{exp}(\mathbf{e}^{\top}_{u}\mathbf{e}_{i}/\tau)}{\sum_{j\in\mathcal{B}_{ab}}\mathrm{exp}(\mathbf{e}^{\top}_{u}\mathbf{e}_{j}/\tau)},$
@@
where $\mathcal{B}_{u}$ and $\mathcal{B}_{i}$ denote the embeddings of $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ within the same batch, $\mathcal{B}_{ui}=[\mathcal{B}_{u},\mathcal{B}_{i}]$ denotes the concatenation of $\mathcal{B}_{u}$ and $\mathcal{B}_{i}$. While this approach appears to be sound, the CL loss disregard the distance between $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, focusing solely on the computation of $\mathbf{e}_{u}$ with $\mathbf{e}_{j}$, resulting in bias in the obtained self-supervised signals.
Figure 4. Distinction in positive pairs and negative pairs samples selection between the high-order contrastive objective and other contrastive objectives.
To enable our method to better aggregate with more negative samples $\mathcal{B}_{ui}$, we present an advanced fusion CL loss that is specifically adapted to high-order contrastive views:
(11) $\mathcal{C}_{fusion}(\mathbf{e}_{u},\mathbf{e}_{i})$ $=-\sum_{<u,i>\in\mathcal{B}}\mathrm{log}\frac{\mathrm{exp}(\mathrm{e}^{\top}_{u}\mathrm{e}_{i}/\tau)}{\sum_{j\in\mathcal{B}_{ui}}\mathrm{exp}((\mathrm{e}^{\top}_{u}\mathrm{e}_{j}+\mathrm{e}^{\top}_{i}\mathrm{e}_{j})/\tau)}.$
@@
By separately calculating the similarity between view $a$ and view $b$ with all negative samples, we effectively fuse the self-supervised signals of different views and maximize the mutual information between positive pairs. Meanwhile, to effectively illustrate how HFGCL combines more negative samples, we show the positive pairs and selected negative samples for various contrastive methods in Fig. 4. Since our high-order contrastive views are derived from two different samples, in the selection of negative samples, unlike traditional contrastive views where only one set can be chosen as negative samples, we can simultaneously use both as negative samples for contrastive loss. Furthermore, our fusion CL loss can effectively aggregate this additional negative sample information and naturally fuse all the self-supervised signals.
@@
Through the study of relationships between positive pairs and negative samples, we have simplified the diverse contrastive objectives traditionally used to capture different self-supervised signals. By utilizing a fusion CL loss, we efficiently obtain the richer self-supervised signals, enabling personalized recommendations.
3.4. Model Optimization
For the purpose of enhancing our contrastive model’s suitability for the primary recommendation task, we implement a multi-task training strategy to optimize the model’s parameters. The overall loss function for HFGCL is defined as follows:
@@ 3.4. Model Optimization
3.5. Theoretical Analysis
Table 2. Cosine similarity of different embeddings.
@@ 3.5. Theoretical Analysis
$\frac{\partial\mathcal{C}_{fusion}}{\partial\mathbf{e}_{u}}$ $=\frac{f}{\tau}\cdot\frac{-\mathbf{e}_{i}+\sum_{j\in\mathcal{B}_{ui}}(\mathbf{e}_{i}+\mathbf{e}_{j})}{\sum_{j\in\mathcal{B}_{ui}}(\mathrm{exp}((\mathbf{e}_{u}+\mathbf{e}_{i})^{\top}\mathbf{e}_{j})/\tau)},$
@@ 4.1. Experimental Settings
Datasets. In selecting datasets to validate the effectiveness of our HFGCL, we choose three datasets that are widely recognized and frequently used for benchmarking in numerous studies: Amazon-book (He et al., 2020), Yelp2018 (Sang et al., 2024), Tmall (Ren et al., 2023). For all datasets, we consider all ratings ‘¿3’ as presence of interactions (i.e., presence of interaction is 1 otherwise 0). We filter out users with less than 10 interactions to ensure the validity of the recommendation. Subsequently, we divide the training and test data into ‘80%’ and ‘20%’ for every users. The details of all the datasets are shown in Table 3.
@@
Baselines. To demonstrate the effectiveness of HFGCL, we select numerous state-of-the-art methods for comparison.
  • MF-based method: BPR-MF (Koren et al., 2009).
  • GCN-based methods: NGCF (Wang et al., 2019), LightGCN (He et al., 2020) and CAGCN*(Wang et al., 2023).
  • GCL-based methods: SGL (Wu et al., 2021), NCL (Lin et al., 2022), SimGCL(Yu et al., 2022b), CGCL(He et al., 2023), VGCL (Yang et al., 2023b), LightGCL (Cai et al., 2023), RecDCL (Zhang et al., 2024a) and BIGCF (Zhang et al., 2024b).
Evaluation Indicators. To evaluate the recommendation efficacy of the HFGCL method, we choose the widely used metrics (Zhang et al., 2023; Yu et al., 2022b), Recall@K and NDCG@K (K=10, 20).
Hyperparameters. To ensure fair and consistent comparisons, all experiments run on a Linux system equipped with a GeForce RTX 2080Ti GPU. We implement HFGCL in the Pytorch environment111. The batch size for all methods are set to 4096, 2048 and 4096 for Amazon-book, Yelp2018 and Tmall datasets, respectively. Except for RecDCL, all methods use an embedding size of 64, while RecDCL uses 2048, as it primarily studies variations across different embedding sizes. Embeddings are initialized using the Xavier strategy (Glorot and Bengio, 2010). Adam is used as the optimizer by default. For methods utilizing the GCN encoder, the number of GCN layers is chosen from {1, 2, 3}. Specifically, for HFGCL, the number of GCN layers $L$ is set to 3, the high-order information $h$ is varied within {1, 2, 3}, and the $\tau$ parameter is selected from {0.20, 0.22, 0.24, 0.26, 0.28, 0.30}. The regularization weights $\lambda_{1}$ is chosen from {0.1,0.5,1.0, 2.5}, and $\lambda_{2}$ is set from {1e-3, 1e-4, 1e-5, 1e-6}.
@@
BPR-MF 0.0170 0.0308 0.0182 0.0239 0.0278 0.0486 0.0317 0.0394 0.0312 0.0547 0.0287 0.0400
NGCF 0.0199 0.0337 0.0200 0.0262 0.0331 0.0579 0.0368 0.0477 0.0374 0.0629 0.0351 0.0465
LightGCN 0.0228 0.0411 0.0241 0.0315 0.0362 0.0639 0.0414 0.0525 0.0435 0.0711 0.0406 0.0530
CAGCN* 0.0300 0.0506 0.0305 0.0400 0.0414 0.0711 0.0470 0.0590 0.0524 0.0840 0.0495 0.0726
DirectAU 0.0296 0.0506 0.0297 0.0401 0.0414 0.0703 0.0477 0.0583 0.0475 0.0752 0.0443 0.0576
GraphAU 0.0300 0.0502 0.0310 0.0400 0.0401 0.0691 0.0463 0.0574 0.0517 0.0840 0.0488 0.0625
SGL 0.0263 0.0478 0.0281 0.0379 0.0395 0.0675 0.0448 0.0555 0.0457 0.0738 0.0434 0.0556
NCL 0.0266 0.0481 0.0284 0.0373 0.0403 0.0685 0.0458 0.0577 0.0459 0.0750 0.0429 0.0553
SimGCL 0.0313 0.0515 0.0334 0.0414 0.0424 0.0721 0.0488 0.0601 0.0559 0.0884 0.0536 0.0674
CGCL 0.0274 0.0483 0.0284 0.0380 0.0404 0.0690 0.0452 0.0560 0.0542 0.0880 0.0510 0.0655
VGCL 0.0312 0.0515 0.0332 0.0410 0.0425 0.0715 0.0485 0.0587 0.0557 0.0880 0.0533 0.0670
LightGCL 0.0303 0.0506 0.0318 0.0397 0.0377 0.0657 0.0437 0.0539 0.0531 0.0832 0.0533 0.0637
RecDCL 0.0311 0.0525 0.0318 0.0407 0.0408 0.0690 0.0464 0.0567 0.0527 0.0853 0.0492 0.0632
BIGCF 0.0294 0.0500 0.0320 0.0398 0.0431 0.0730 0.0497 0.0603 0.0547 0.0876 0.0524 0.0664
@@
Improv.10.54% 7.81% 12.57% 10.63% 3.94% 3.01% 3.62% 3.83% 3.22% 3.05% 3.54% 3.41%

### v4

@@ 4. HFGCL: HIGH-ORDER FUSION GRAPH CONTRASTIVE LEARNING
In this section, we propose a High-order Fusion Graph Contrastive Learning (HFGCL) framework in Fig. 2. HFGCL leverages high-order information to generate contrastive views and introduces a fusion CL loss, effectively integrating self-supervised signals from multi-scale CL objectives to alleviate the data sparsity issue.
4.1. User-Item Contrastive Learning Paradigm
Following the analysis in Section 3, we urgently need to find a novel contrastive learning paradigm. Inspired by the natural language processing (NLP) (Gao et al., 2022b), which generates contrastive views by applying the dropout function to input sentence. We believe that similar contrastive views also exist in graph-based recommendation tasks.
Considering a different perspective, we directly treat users and items as contrastive views. Given that users and items are first-order neighbors of each other, their high-order information are essentially identical. As graph convolution (GCN) progresses, users and items gradually capture neighborhood information, resulting in their embeddings becoming progressively closer. Therefore, GCN process ultimately generates a set of distinctive contrastive views.
The previous contrastive views originate from the same sample ($e.g.,$ $\mathbf{e}_{u}^{\prime}$ and $\mathbf{e}_{u}^{\prime\prime}$ from $\mathbf{e}_{u}$, as shown in Eq. 4), necessitating that a set of views can be selected as negative samples ($e.g.,$ $\sum_{i\in B}\mathbf{e}_{u_{i}}^{\prime\prime}$):
@@ 4.2. High-order Contrastive Views
Many studies (Wang et al., 2019) use aggregation functions to integrate layer-wise embeddings into the final encoder. Though the encoder effectively aggregates information from both the sample itself and its neighborhoods, we find that the low-order information is extremely unique. This uniqueness can seriously affect the similarity between positive pairs, thereby interfering with the mutual information of them.
@@
4.3. Multi-scale Fusion of Self-supervised Signals
In recommender systems, data augmentation typically involve augmenting the original user or item embeddings with multiple GCN and model modeling operations to generate different contrastive views. When calculating CL loss, different views of the same sample are regarded as positive pairs, while a set of augmented views are selected as negative pairs. Among them, it is not difficult to find that the negative pairs used in the previous CL loss come from a set of augmented views. In contrast, for the high-order contrastive views propose in Section 4.2, we give the corresponding CL loss in Eq. 6. It is differently from the previous negative pairs sampling approach, the negative pairs come from the opposing views. This difference leads to the self-supervised signals of users and items themselves are missing.
Original CL Loss. To be able to obtain this part of the self-supervised signal, we attempt to treat the samples themselves as negative pairs and introduce the corresponding CL loss:
@@ 4.3. Multi-scale Fusion of Self-supervised Signals
where $a$ and $b$ denote different users/items samples. Unlike previous GCL-based methods (Yu et al., 2022; Yang et al., 2023b), the positive pairs ($i.e.,$ $\langle a,a\rangle$) of samples are identical. Subsequently, we combine the CL loss associated with the users and the items themselves:
@@
After obtaining the CL objectives for both users and items themselves, we combine these objectives with main CL loss (Eq. 6):
@@
Fusion CL Loss. Most studies obtain self-supervised signals through traditional CL objectives ($e.g.,$ separately obtaining signals from users and items). Although effective, we argue that differences in these self-supervised signals lead to suboptimal recommendations. Instead, in our CL objective, we propose four CL objectives to capture self-supervised signals. While proposed CL objectives capture more signals, in fact, different self-supervised signals make them extremely difficult to effectively integrate these self-supervised signals. It motivates us to further improve our CL objective to fuse these diverse self-supervised signals. Considering that, unlike the traditional CL loss, our proposed CL loss allows both users and items to serve as negative pairs. A simple strategy is to directly treat the concatenation of two contrastive views as negative pairs when constructing contrastive objectives:
(12) $\mathcal{C}_{con}(\mathbf{e}_{u},\mathbf{e}_{i})$ $=-\sum_{\langle u,i\rangle\in\mathcal{B}}\mathrm{log}\frac{\mathrm{exp}(\mathbf{e}^{\top}_{u}\mathbf{e}_{i}/\tau)}{\sum_{j\in\mathcal{B}_{ui}}\mathrm{exp}(\mathbf{e}^{\top}_{u}\mathbf{e}_{j}/\tau)},$
@@
where $\mathcal{B}_{u}$ and $\mathcal{B}_{i}$ denote the embeddings of $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ within the same batch, $\mathcal{B}_{ui}=[\mathcal{B}_{u},\mathcal{B}_{i}]$ denotes the concatenation of $\mathcal{B}_{u}$ and $\mathcal{B}_{i}$. While this strategy appears to be sound, the CL loss disregard the distance between $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, focusing solely on the computation of $\mathbf{e}_{u}$ with $\mathbf{e}_{j}$, resulting in bias in the obtained self-supervised signals.
Figure 3. HFGCL Optimization Process. This process optimizes the primary objective, while fusing user, item, and user-item self-supervised signals from diverse contrastive views via a fusion CL objective.
To eliminate such biases, we propose a fused embedding representation $\mathbf{e}^{*}_{ui}$. Specifically, we dynamically reconstruct and fuse the high-order neighborhood information of users and items:
(13) $\mathbf{e}^{*}_{ui}=2\sum_{l=h}^{L}\alpha\sum_{i\in\mathcal{Z}_{u}}\frac{1}{\sqrt{|\mathcal{Z}_{u}|}\sqrt{|\mathcal{Z}|_{i}}}\mathbf{e}_{u}^{(l)}+(1-\alpha)\sum_{u\in\mathcal{Z}_{i}}\frac{1}{\sqrt{|\mathcal{Z}_{u}|}\sqrt{|\mathcal{Z}_{i}|}}\mathbf{e}_{i}^{(l)},$
where $\mathcal{Z}_{u}$ and $\mathcal{Z}_{i}$ denote the neighbors of users and items, respectively, $\alpha$ represents the parameter to adjust the contribution of user and item embeddings in the fusion process, $\mathbf{e}_{u}^{(h)}$ and $\mathbf{e}_{i}^{(h)}$ denote the embeddings obtained by stacking multiple GCN layers on the initial embeddings $\mathbf{e}_{u}^{(0)}$ and $\mathbf{e}_{u}^{(0)}$, respectively. Subsequently, to better integrate the contrastive objectives generated by user, item, and user-item contrastive views, we present an advanced fusion CL loss that is specifically adapted to high-order contrastive views:
(14) $\mathcal{C}_{fusion}(\mathbf{e}_{u},\mathbf{e}_{i})$ $=-\sum_{\langle u,i\rangle\in\mathcal{B}}\mathrm{log}\frac{\mathrm{exp}(\mathbf{e}^{\top}_{u}\mathbf{e}_{i}/\tau)}{\sum_{j\in\mathcal{B}_{ui}}\mathrm{exp}(\mathbf{e}^{*\top}_{ui}\mathbf{e}_{j})/\tau)},$
@@
By calculating the similarity between fused embedding $\mathbf{e}^{*}_{ui}$ and richer negative pairs $\mathcal{B}_{ui}$, we can effectively integrate the self-supervised signals from diverse views and maximize the mutual information between the positive pair $\langle u,i\rangle$. Meanwhile, to effectively illustrate how HFGCL optimizes the model, we provide the training process in Fig. 3 and demonstrate the signal fusion process. Since our high-order contrastive views select user-item interactions as positive pairs, unlike traditional contrastive methods that can only choose one set as negative pairs, we can simultaneously treat two sets of samples as negative pairs $\mathcal{B}_{ui}$ for contrastive learning. Furthermore, our fusion CL loss can effectively aggregate this additional negative pairs information and naturally fuse all the self-supervised signals.
@@
Through studying the relationships between multi-scale contrastive objectives, we simplify the various contrastive objectives traditionally used to capture different self-supervised signals. By utilizing a fusion CL loss, we efficiently obtain the richer self-supervised signals, enabling personalized recommendations.
4.4. Method Optimization
For the purpose of enhancing our contrastive method’s suitability for the primary recommendation task, we implement a multi-task training strategy to optimize the method’s parameters. The overall loss function for HFGCL is defined as follows:
@@ 4.4. Method Optimization
4.5. Method Analysis
4.5.1. Time Complexity Analysis
To demonstrate the efficiency of our HFGCL, we investigate the time complexity of the GCN- and GCL-based methods. Specifically, the embeddings in HFGCL are generated through graph convolution, yielding a time complexity of $\mathcal{O}(2Ld|\mathcal{E}|)$, where $|\mathcal{E}|$ denotes the number of edge, $d$ denotes the embedding dimensions, $L$ denotes the GCN layer. Additionally, since calculating the CL loss requires fusing the self-supervised signals from users, items and user-item contrastive views, the time complexity becomes $\mathcal{O}(\mathcal{B}d(1+4\mathcal{B}))$.
Table 3. Time complexity of proposed HFGCL method and the GCN- and CL-based baselines.
Method Encoding CL Loss
NGCF (Wang et al., 2019) $\mathcal{O}(2Ld(|\mathcal{E}|+|\mathcal{V}|))$
LightGCN (He et al., 2020) $\mathcal{O}((2Ld|\mathcal{E}|)$
SGL-ED (Wu et al., 2021) $\mathcal{O}(2(1+2\rho)Ld|\mathcal{E}|)$ $\mathcal{O}(2\mathcal{B}d(1+\mathcal{B}))$
SimGCL (Yu et al., 2022) $\mathcal{O}(6Ld|\mathcal{E}|)$ $\mathcal{O}(2\mathcal{B}d(1+\mathcal{B}))$
NCL (Lin et al., 2022) $\mathcal{O}(2Ld|\mathcal{E}|+d|\mathcal{K}||\mathcal{V}|)$ $\mathcal{O}(4\mathcal{B}d+2\mathcal{B}|\mathcal{V}|))$
CGCL (Yang et al., 2023b) $\mathcal{O}(2Ld|\mathcal{E}|)$ $\mathcal{O}(6\mathcal{B}d+3\mathcal{B}|\mathcal{V}|)$
VGCL (Yang et al., 2023b) $\mathcal{O}(2Ld|\mathcal{E}|+d|\mathcal{K}||\mathcal{V}|)$ $\mathcal{O}(4\mathcal{B}d(1+\mathcal{B}))$
LightGCL (Cai et al., 2023) $\mathcal{O}(2Ld(|\mathcal{E}|+q|\mathcal{V}|))$ $\mathcal{O}(2\mathcal{B}d+\mathcal{B}|\mathcal{V}|)$
BIGCF (Zhang et al., 2024b) $\mathcal{O}(2Ld|\mathcal{E}|+d|\mathcal{K}||\mathcal{V}|)$ $\mathcal{O}(5\mathcal{B}d(1+\mathcal{B}))$
HFGCL(Ours) $\mathcal{O}(2Ld|\mathcal{E}|)$ $\mathcal{O}(\mathcal{B}d(1+4\mathcal{B}))$
Table 3 summarizes the time complexity of HFGCL and several classical baselines, where $\rho$ denotes the probability of retaining edges or nodes in SGL-ED, $|\mathcal{K}|$ represents the number of collective intent nodes in BIGCF and the number of clusters in the clustering algorithms used for NCL and VGCL, $|\mathcal{V}|=M+N$ is the total number of user and item nodes, and $q$ indicates the required rank for SVD in LightGCL. Methods based on GCL generate contrastive views through data augmentation ($e.g.,$ SGL-ED and SimGCL), which require additional graph convolution operations and significantly increase the overall time complexity. Even methods that avoid these extra convolutions often treat all nodes as negative samples to maximize the mutual information of positive pairs ($e.g.,$ NCL and CGCL), while some approaches involve time-consuming specialized modeling processes, such as graph reconstruction in VGCL and intent modeling in BIGCF. In contrast, HFGCL eliminates the need for additional data augmentation or specialized modeling operations during training. By introducing high-order contrastive views and proposing a fused contrastive loss, HFGCL achieves efficient contrastive learning, significantly enhancing both efficiency and recommendation performance.
4.5.2. Theoretical Analysis
Table 4. Comparison of user-item pair Cosine Similarities under different GCN layers on three datasets.
@@ 4.5.2. Theoretical Analysis
$\frac{\partial\mathcal{C}_{fusion}}{\partial\mathbf{e}_{u}}$ $=\frac{f}{\tau}\cdot\frac{-\mathbf{e}_{i}+\sum_{j\in\mathcal{B}_{ui}}(\mathbf{e}_{i}+\mathbf{e}_{j})}{\sum_{j\in\mathcal{B}_{ui}}(\mathrm{exp}(\mathbf{e}^{{*\top}}_{ui}\mathbf{e}_{j})/\tau)},$
@@ 5.1. Experimental Settings
Datasets. In selecting datasets to validate the effectiveness of our HFGCL, we choose three datasets that are widely recognized and frequently used for benchmarking in numerous studies: Amazon-book (He et al., 2020), Yelp2018 (Sang et al., 2025), Tmall (Ren et al., 2023). For all datasets, we consider all ratings ‘¿3’ as presence of interactions (i.e., presence of interaction is 1 otherwise 0). We filter out users with less than 10 interactions to ensure the validity of the recommendation. The details of all the datasets are shown in Table 5.
@@
BPR-MF (UAI’2009) 0.0170 0.0308 0.0182 0.0239 0.0278 0.0486 0.0317 0.0394 0.0312 0.0547 0.0287 0.0400
NGCF (SIGIR’2019) 0.0199 0.0337 0.0200 0.0262 0.0331 0.0579 0.0368 0.0477 0.0374 0.0629 0.0351 0.0465
LightGCN (SIGIR’2020) 0.0228 0.0411 0.0241 0.0315 0.0362 0.0639 0.0414 0.0525 0.0435 0.0711 0.0406 0.0530
DirectAU (KDD’2022) 0.0296 0.0506 0.0297 0.0401 0.0414 0.0703 0.0477 0.0583 0.0475 0.0752 0.0443 0.0576
GraphAU (WSDM’2023) 0.0300 0.0502 0.0310 0.0400 0.0401 0.0691 0.0463 0.0574 0.0517 0.0840 0.0488 0.0625
Mult-VAE (WWW’2018) 0.0224 0.0407 0.0239 0.0315 0.0335 0.0584 0.0359 0.0450 0.0467 0.0740 0.0423 0.0552
CVGA (TOIS’2023) 0.0290 0.0492 0.0302 0.0379 0.0407 0.0694 0.0467 0.0571 0.0540 0.0854 0.0517 0.0648
DiffRec (SIGIR’2023) 0.0310 0.0514 0.0333 0.0418 0.391 0.0665 0.0447 0.0556 0.0485 0.0792 0.0473 0.0612
SGL-ED (SIGIR’2021) 0.0263 0.0478 0.0281 0.0379 0.0395 0.0675 0.0448 0.0555 0.0457 0.0738 0.0434 0.0556
SimGCL (SIGIR’2022) 0.0313 0.0515 0.0334 0.0414 0.0424 0.0721 0.0488 0.0601 0.0559 0.0884 0.0536 0.0674
NCL (WWW’2022) 0.0266 0.0481 0.0284 0.0373 0.0403 0.0685 0.0458 0.0577 0.0459 0.0750 0.0429 0.0553
CGCL (SIGIR’2023) 0.0274 0.0483 0.0284 0.0380 0.0404 0.0690 0.0452 0.0560 0.0542 0.0880 0.0510 0.0655
VGCL (SIGIR’2023) 0.0312 0.0515 0.0332 0.0410 0.0425 0.0715 0.0485 0.0587 0.0557 0.0880 0.0533 0.0670
LightGCL (ICLR’2023) 0.0303 0.0506 0.0318 0.0397 0.0377 0.0657 0.0437 0.0539 0.0531 0.0832 0.0533 0.0637
RecDCL (WWW’2024) 0.0311 0.0525 0.0318 0.0407 0.0408 0.0690 0.0464 0.0567 0.0527 0.0853 0.0492 0.0632
BIGCF (SIGIR’2024) 0.0294 0.0500 0.0320 0.0398 0.0431 0.0730 0.0497 0.0603 0.0547 0.0876 0.0524 0.0664
SCCF (KDD’2024) 0.0287 0.0491 0.0294 0.0399 0.0423 0.0718 0.0489 0.0595 0.0478 0.0772 0.0453 0.0580
@@
Improv.10.54% 7.81% 12.57% 9.57% 3.94% 3.01% 3.62% 3.83% 3.22% 3.05% 3.54% 3.41%
Baselines. To demonstrate the effectiveness of HFGCL, we select numerous state-of-the-art baselines for comparison.
  • Only MF-based method: BPR-MF (Koren et al., 2009).
  • GCN-based methods: NGCF (Wang et al., 2019) and LightGCN (He et al., 2020).
  • Generative-based methods: Mult-VAE (Liang et al., 2018), CVGA (Zhang et al., 2023) and DiffRec (Wang et al., 2023).
  • CL-based methods: SGL-ED (Wu et al., 2021), NCL (Lin et al., 2022), SimGCL(Yu et al., 2022), CGCL (He et al., 2023), VGCL (Yang et al., 2023b), LightGCL (Cai et al., 2023), RecDCL (Zhang et al., 2024a), BIGCF (Zhang et al., 2024b) and SCCF (Wu et al., 2024).
Evaluation Indicators. To evaluate the recommendation efficacy of the HFGCL method, we choose the widely used metrics (Zhang et al., 2023; Yu et al., 2022), Recall@K and NDCG@K (K=10, 20).
Hyperparameters. To ensure fair and consistent comparisons, all experiments run on a Linux system equipped with a GeForce RTX 2080Ti GPU. We implement HFGCL in the Pytorch environment [^1] https://anonymous.4open.science/r/HFGCL-main-2F79. The batch size for all methods are set to 4096, 2048 and 4096 for Amazon-book, Yelp2018 and Tmall datasets, respectively. Except for RecDCL, all methods use an embedding size of 64, while RecDCL uses 2048, as it primarily studies variations across different embedding sizes. Embeddings are initialized using the Xavier strategy (Glorot and Bengio, 2010). Adam is used as the optimizer by default. For methods utilizing the GCN encoder, the number of GCN layers is chosen from {1, 2, 3}. Specifically, for HFGCL, the number of GCN layers $L$ is set to 3, the high-order information $h$ is varied within {1, 2, 3}, and the $\tau$ parameter is selected from {0.20, 0.22, 0.24, 0.26, 0.28, 0.30}. The regularization weights $\lambda_{1}$ is chosen from {0.1,0.5,1.0, 2.5}, $\lambda_{2}$ is set from {1e-3, 1e-4, 1e-5, 1e-6}, and $\alpha$ is generally set to 0.5 by default.